OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9

OpenStudy (nurali):

Start with the slope formula. \[m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] Plug in y2=9, y1=6, x2=2 and x1=1 \[m=\frac{ 9-6 }{ 2-1 }\] \[m=3\] So the slope of the line that goes through the points (1, 6) and (2, 9) is m=3 Now let's use the point slope formula: Start with the point slope formula \[y-y _{1}=m(x-x _{1})\] Plug in m=3, x1=1 and y1=6 \[y-6=3(x-1)\] \[y-6=3x-3\] \[y=3x-3+6\] \[y=3x+3\]